5x + y = 9
10x - 7y = -18
Solving systems of Equations by Elimination
Don't understand how to do can you show your work please.
Thank you

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

5x + y = 9
10x - 7y = -18
----------
10x + 2y = 18 1st eqn times 2
10x - 7y = -18
------------------- Subtract
9y = 36
y = 4
------------
5x + y = 9
5x + 4 = 9
5x = 5
x = 1

Answer by ikleyn(52778)

.
On Elimination method solving two linear equations in two unknowns see the lesson

      Solution of the linear system of two equations in two unknowns by the Elimination method

in this site.